Assume you bought a share of a stock for $20 one year ago, sold it for $25 today. During the year, you received $2 as dividends. What is your capital gain? What is you dividend yield? What is your total return?

1.

25%, 10%, 35%

2.

25%, 10%, 30%

3.

20%, 10%, 30%

4.

15%, 10%, 25%
Answer #1

HI

Here bought share price = $20

sold share price = $25

So capital gain = (sold share price - bought share price)/bought share price

=25-20/20 = 5/20 = 25%

dividend yield = dividend / bought share price

=2/20 = 10%

total return = dividend + profit on share bought

=2+25-20 = $7

total return% = return/share bought =7/20 = 35%

so option first is correct here

Thanks